2.0 Define the electronvolt.

(2)

oint A. -9 +4x10 C

Β 2.1 In the Millikan experiment, an oil drop carries four electronic charges (3)

and has a mass of 1.8 x 10-12 g. It is held almost at rest between two horizontal charged plates 1.8 cm apart. What voltage must there be between the two charged plates? (Take electron charge e = 1.60 x 10-19 C)

3.0 Compute the energy stored in a 60𝑝𝐹 capacitor (5)

a) When Charged to a potential difference of 2π‘˜π‘£

b) When the charge on each plate is 30𝑛𝐢


1
Expert's answer
2022-03-26T15:32:01-0400

"E=\frac Fq=\frac{mg}q-\frac{kq}{r^2}=27 .6~kV."

"W_1=\frac{CU^2}2=120~\mu J,"

"W_2=\frac{q^2}{2C}=7.5~\mu J."